OpenStudy (revircs):

Need help with implicit differentiation. Consider the implicit function y*(xy+1)=4(x+4) a) what is the implicit derrivative? b)Find the tangent line passing through the point (1,4) c) Does the graph have any horizontal lines? if so, where? Any guidance on this problem would be greatly appreciated.

OpenStudy (freckles):

have you tried differentiating both sides?

OpenStudy (solomonzelman):

You are applying the same rules for taking the derivative ! (Every time when you differentiate something that contains y you will multiply times y\('\) -- this is because y is really a function of x, and therefore it gets its own chain-rule, JUST AS a derivative of sin(4x), \(\sqrt{x^3+4x-3}\), or many other functions would get.)

OpenStudy (revircs):

So I would just need to distribute, then find the derivative of the entire problem? xy^2+y=4x+16 to 2xy * y' + 1*y'=4? then to 2y'=4-2xy to y'=(4-2xy)/2?

OpenStudy (revircs):

@SolomonZelman

OpenStudy (freckles):

hi

OpenStudy (freckles):

\[xy^2+y=4x+16 \\ \text{ you need to use product rule for the } (xy^2)' \text{ part }\]

OpenStudy (freckles):

\[\frac{d}{dx}(xy^2)= y^2 \cdot \frac{d}{dx}(x)+x \cdot \frac{d}{dx}(y^2) \\ =?\]

OpenStudy (revircs):

so after that, I found \[\frac{ dy }{ dx }=\frac{ 4-y^2 }{ 4xy },\] But I feel like I may have done something wrong.

OpenStudy (revircs):

Nevermind, I found my mistake. The derivative should actually be \[\frac{ dy }{ dx }=\frac{ 4-y^2 }{ 2xy+1 }\]

OpenStudy (revircs):

Then to find the tangent line at 1,4 I believe we would need to plug these x and y values into the derivative to give us our slope, so: \[\frac{ 4-(4)^2 }{ 2(1)(4)+1 } = \frac{ -12 }{ 9}\]

OpenStudy (revircs):

Using that in the formula y-y_1=m(x-x_1) would give us \[y=\frac{ -4 }{ 3 }(x-1)+4\]

OpenStudy (revircs):

So, I just need help finding if it has any horizontal tangents, and where they are located. I believe we would need to set our derivative equal to zero, but I dont know what to do now that there are an x and a y.

OpenStudy (freckles):

hey

OpenStudy (revircs):

Hello

OpenStudy (freckles):

\[xy^2+y=4x+16 \\ x \cdot 2y y'+y^2+y'=4 \\ y'(2xy+1)=4-y^2 \\ y'=\frac{4-y^2}{2xy+1}\] your y' looks awesome!

OpenStudy (freckles):

\[y'|_{(1,4)}=\frac{4-4^2}{2(1)(4)+1}=\frac{-12}{9}=\frac{-4}{3} \\ y-f(1)=f'(1)(x-1) \\ y-4=\frac{-4}{3}(x-1)\] your tangent line at (1,4) looks great so now you are looking for horizontal tangents you say

OpenStudy (freckles):

horizontal tangents occur when y'=0

OpenStudy (freckles):

y' we have written as a fraction fractions are 0 when the numerator is 0

OpenStudy (freckles):

\[xy^2+y=4x+16 \\ x \cdot 2y y'+y^2+y'=4 \\ y'(2xy+1)=4-y^2 \\ y'=\frac{4-y^2}{2xy+1} \\ 4-y^2=0\]

OpenStudy (freckles):

4-y^2=0 this equations is going to help us find the horizontal tangents

OpenStudy (revircs):

okay :)

OpenStudy (freckles):

notice this equation is just in terms of y so we solve for y and find x by pluggin into the original equation

OpenStudy (revircs):

So y would be y= +/- 2?

OpenStudy (freckles):

right now plug into original function to find the corresponding x values

OpenStudy (freckles):

hey and also when we find these x values...

OpenStudy (freckles):

we want to make sure it doesn't make y' undefined when pluggin in the point like we do not want the 1+2xy to be 0

OpenStudy (freckles):

anyways let's see what happens

OpenStudy (revircs):

\[x(2)^2+(2)=4x+16\] \[4x+2=4x+16\] since these can't be equal, could we assume that there are no tangent lines?

OpenStudy (freckles):

oops I replace the x with 2

OpenStudy (revircs):

I thought that the y was 2?

OpenStudy (freckles):

yeah that is why I said oops :) \[xy^2+y=4x+16 \\ y=2 \\ 4x+4=4x+16 \\ y=-2 \\ 4x-2=4x+16 \] there is no x for either of trhese

OpenStudy (freckles):

so there are no horizontal tangents

OpenStudy (revircs):

ooh, haha I thought you wanted me to put x as 2, sorry about that!

OpenStudy (revircs):

Okay, thank you so much!

OpenStudy (freckles):

nope it was my oops not your oops :)

OpenStudy (revircs):

So to find this tangent line, we took the top part of the derivative. If we wanted to find vertical lines, would we set the bottom equal to 0?

OpenStudy (freckles):

yep vertical lines can happen when the denominator of y' is 0

OpenStudy (revircs):

ooh, so having the top of the fraction equal to 0 would make the entire thing 0, but with a 0 in the denominator, it would be undefined, making it vartical.

OpenStudy (revircs):

Vertical*

OpenStudy (freckles):

well it could make it vertical

OpenStudy (freckles):

it could be just no continuous there and that is y' doesn't exist

OpenStudy (revircs):

oh yeah, thats a good point

OpenStudy (freckles):

\[xy^2+y=4x+16 \\ y'=\frac{4-y^2}{2xy+1} \\ 2xy+1 =0 \\ xy=\frac{-1}{2} \\\\ xy^2+y=4x+16 \\ \text{ multiply } x \text{ on both sides } \\ x^2y^2+xy=4x^2+16x \\ \text{ replace } xy \text{ with } \frac{-1}{2} \\ (\frac{-1}{2})^2+\frac{-1}{2}=4x^2+16x \\ \] \[\frac{1}{4}-\frac{1}{2}=4x^2+16 x \\ 1-2=16x^2+64x \\ 16x^2+64x+1=0\]

OpenStudy (revircs):

What does that final part tell us?

OpenStudy (freckles):

\[x=-2 \pm \frac{3 \sqrt{7}}{4} \\ x=-2+\frac{3 \sqrt{7}}{4} \approx -0.0157 \\ x=-2-\frac{3 \sqrt{7}}{4} \approx -3.984\] to see if this point actually exist we would have to plug in this x into original to see if there is a corresponding y that exists http://www.wolframalpha.com/input/?i=%28-2%2B3+sqrt%287%29%2F4%29y%5E2%2By%3D4%28-2%2B3+sqrt%287%29%2F4%29%2B16 \[\text{ this above link says when } x=-2+\frac{3 \sqrt{7}}{4} \text{ that we have } y=16+6 \sqrt{7} \approx 31.875 \\ \\\] http://www.wolframalpha.com/input/?i=%28-2-3+sqrt%287%29%2F4%29y%5E2%2By%3D4%28-2-3+sqrt%287%29%2F4%29%2B16 \[\text{ this above link says when } x=-2+\frac{3 \sqrt{7}}{4} \text{ that we have } y=16-6 \sqrt{7} \approx 0.12549 \\ \\\] hmm... I wonder if xy=-1/2...

OpenStudy (freckles):

yep just verified that last sentence

OpenStudy (revircs):

Since theres an actual point in both, wold this imply no vertical slopes either?

OpenStudy (freckles):

yep we have horizontal tangents when the y' numerator is 0 but its denominator isn't 0 we have vertical tangents when y' denominator is 0 but its numerator isn't 0

OpenStudy (revircs):

Awesome, thank you!

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=+xy%5E2%2By%3D4x%2B16 I don't get wolfram's graph

OpenStudy (freckles):

If I zoom maybe I see a vertical tangent close to (-0.015,31) http://www.wolframalpha.com/input/?i=+xy%5E2%2By%3D4x%2B16%2Cx+in+%28-1%2C0%29

OpenStudy (freckles):

I just realized I made a type-o above \[\text{ this above link says when } x=-2-\frac{3 \sqrt{7}}{4} \text{ that we have } y=16-6 \sqrt{7} \approx 0.12549 \\ \\\]

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=+xy%5E2%2By%3D4x%2B16%2Cx+in+%28-4%2C-3%29%2Cy+in+%280%2C.5%29 for the one that is close to (-3.98,.12) I guess I just had to do a lot of zooming in

OpenStudy (revircs):

@freckles Even though there was no x values that made 4x+2=4x+16?

OpenStudy (freckles):

what is this question about?

OpenStudy (revircs):

The graph you put up showed that there was a vertical tangent line at (-3.98,.12), but when we tried to solve it algebraically, it showed that there wasnt.

OpenStudy (freckles):

no

OpenStudy (freckles):

I think you are getting horizontal and vertical mixed up

OpenStudy (freckles):

we had no horizontal tangents because 4x+4=4x+16 or 4x-2=4x+16 didn't have a solution but we did find two vertical tangents

OpenStudy (revircs):

oooh, yeah, never mind